# Vertex-based Graph Neural Solver and its Application to Linear Elasticity Equations

### v1

@@ Abstract
The numerical solution of linear elasticity equations on unstructured grids presents significant difficulties, especially in the presence of heterogeneous and anisotropic materials. To tackle these challenges, we introduce a novel vertex-based graph Fourier neural solver. This framework seamlessly integrates a block-Jacobi smoothing stage, known for rapidly attenuating certain high-frequency errors, with a learnable Fourier-domain module designed to correct errors across the remained frequency spectrum. Extensive numerical experiments on 2D and 3D problems verify that our proposed solver exhibits superior robustness and efficiency, significantly surpassing the classical smoothed aggregation algebraic multigrid method. Moreover, frequency domain analysis empirically confirms the essential complementarity between the smoother and the neural corrector.
@@
journal: Computer Methods in Applied Mechanics and Engineering
@@ 1 Introduction
The linear elasticity model constitutes the cornerstone of solid mechanics [1, 2, 3], characterizing the stress-strain behavior of deformable bodies under external loads. This framework is ubiquitous across engineering disciplines, ranging from civil and aerospace to mechanical engineering and materials science. Mathematically, the governing equations form a prototypical system of vector-valued partial differential equations (PDEs), which are distinguished by the inherent coupling between vector components. Furthermore, driven by the escalating demand for high-fidelity simulations, solving these equations on complex geometries involving heterogeneous materials imposes stringent requirements on the efficiency and robustness of computational algorithms [4, 5].
@@
The Finite Element Method (FEM) has established itself as the dominant discretization technique, offering exceptional flexibility in handling complex boundaries and multi-physics coupling [6]. Following finite element discretization, the resulting equations typically manifest as large-scale sparse linear systems. To solve such systems, classical iterative algorithms [7]—such as Jacobi, Gauss-Seidel, Conjugate Gradient (CG), and Generalized Minimal Residual (GMRES) iterative methods—combined with preconditioning techniques such as multigrid (MG) [8, 9] or Domain Decomposition methods, have demonstrated good performance. However, when the system exhibits strong heterogeneity, strong anisotropy, or is discretized on complex unstructured mesh topologies, these traditional methods often suffer from significant degradation in convergence rates. Furthermore, the construction of effective preconditioners in such scenarios is non-trivial and often relies on parameter tuning. Consequently, achieving efficient solutions for complex systems while maintaining theoretical convergence guarantees remains a critical challenge in the numerical analysis of PDEs.
@@
In recent years, deep learning technology has made significant inroads into scientific computing, offering novel perspectives on the development of numerical methods. Research on solving physical problems such as linear elasticity has primarily proceeded along two trajectories. The first trajectory utilizes use deep neural network to approximate the solution, such as physics-informed neural networks, which solve specific boundary value problems by embedding PDE residuals into the loss function [10]. The second trajectory focuses on Neural Operators, such as DeepONet [11] and the Fourier Neural Operator (FNO) [12], which aim to learn the mapping from parameter space (e.g., material properties, external forces) to solution space, demonstrating potential as surrogates for traditional numerical solvers [13].
Despite their promise, methods relying entirely on neural networks for surrogate modeling often struggle to meet the high standards of accuracy and numerical stability demanded by engineering computations. This limitation has prompted researchers to shift towards integrating deep learning with traditional iterative methods, establishing a new paradigm known as “Neural Solvers.” Relevant research primarily falls into two categories:
On the one hand, researchers aim to learn key parameters within traditional iterative methods. These approaches preserve the mathematical structure of the original algorithms while utilizing neural networks to optimize specific parameters, thereby typically retaining the theoretical convergence guarantees of the underlying numerical methods. Specific examples include learning optimal weight parameters in Chebyshev iteration [14, 15] or optimal diagonal preconditioners in weighted Jacobi methods [16]; in multigrid methods, learning smoother [17, 18, 19] or prolongation/restriction operators for coarse grids [20, 21, 22]; in domain decomposition methods, learning transmission conditions between subdomains [23] or coarse space constructions [24, 25]; and predicting high-quality initial guesses for Krylov subspace methods [26, 27].
On the other hand, researchers attempt to replace part or all of the preconditioner’s functionality with neural networks. This can be further categorized into two strategies. The first is constructing Deep Learning-based Hybrid Iterative Methods (DL-HIM), where simple iterative methods (such as Jacobi) are used to eliminate high-frequency errors, while neural networks are trained to correct low-frequency errors, forming a frequency-complementary solution strategy [28, 29, 30, 31, 32, 33]. The second strategy involves developing neural operator preconditioning techniques, where neural operators (e.g., GNNs, FNOs) directly learn an approximate inverse of the system matrix, which is then embedded as a preconditioner within Krylov subspace methods [34, 35, 36, 37].
@@
However, existing research on neural solvers has largely concentrated on scalar PDEs or numerical solutions on structured grids. There is a relative scarcity of studies addressing vector-valued PDEs (such as linear elasticity equations) on unstructured grids, which are more prevalent in complex engineering analysis. From the perspective of numerical linear algebra, traditional iterative methods on unstructured grids face multiple hurdles, including convergence slowdowns due to mesh irregularity, ill-conditioning arising from high-contrast coefficients, and the difficulty of automating and generalizing algorithm design. Therefore, constructing a solver framework on unstructured grids that retains the numerical stability of traditional methods while leveraging the adaptability of neural networks remains a significant research challenge.
@@
In light of this, this paper focuses on the efficient numerical solution of linear elasticity problems discretized on unstructured grids. We propose a hybrid solver that integrates block weighted Jacobi smoothing with a Graph Neural Network (GNN)-enhanced Fourier error correction mechanism. GNNs are naturally suited for processing graph-structured data and can directly model the topological connectivity and nodal physical attributes of unstructured meshes, thereby effectively preserving the coupling relationships between displacement components and local interaction mechanisms. The block weighted Jacobi method rapidly attenuates high-frequency error components, which, when combined with frequency-domain error correction, allows for the effective treatment of error modes across the frequency spectrum during the iteration process. The main innovations and contributions of this paper are summarized as follows:
@@
    Building upon the Fourier Neural Solver (FNS) established in [30], we propose a vertex-based, topology-aware iterative neural solver suitable for vector-valued PDEs on unstructured grids. By introducing a Fourier basis transformation network and integrating GNN, this work substantially extends the applicability and scope of the original FNS.
@@
    Through a series of numerical experiments on 2D and 3D linear elasticity problems involving isotropic heterogeneous materials and anisotropic materials, we demonstrate that the proposed solver exhibits superior robustness and solution efficiency. In particular, its performance significantly outperforms the Smoothed Aggregation Algebraic Multigrid (SA-AMG) method [38].
@@
    Based on frequency domain error analysis, we empirically validate the complementarity between the simple iterative operator $\mathcal{B}$ (smoother) and the neural correction operator $\mathcal{H}$ (corrector) in handling error components of different frequencies within the new Neural Solver framework.
@@
The remainder of this paper is organized as follows: Section 2 introduces the necessary theoretical foundations, including the finite element discretization of linear elasticity equations, the implementation of block weighted Jacobi methods using graph structures, and the basic principles of the FNS. Section 3 describes the datasets used for training and testing. Section 4 presents three progressively improved FNS architectures designed to effectively handle vector-valued variable-coefficient PDEs. Finally, Section 5 summarizes the paper and discusses directions for future research.
2 Problem Settings and Preliminaries
In this section, we introduce the mathematical formulations and preliminary concepts essential for the proposed method. We first present the governing equations of the static linear elasticity problem and its discretization using the FEM. Subsequently, we describe the graph representation of such systems to facilitate the application of GNN, specifically detailing the implementation of block matrix-vector multiplication via message passing. Finally, we review the framework of DL-HIM and the standard FNS, discussing their theoretical basis and identifying the limitations that necessitate the generalized approach developed in this work.
2.1 Linear Elasticity Model and Finite Element Discretization
@@
Consider a static linear elasticity system of partial differential equations defined on a bounded domain $\Omega\subset\mathbb{R}^{d}$ (where $d=2,3$). The objective is to find the displacement field $\boldsymbol{u}:\Omega\to\mathbb{R}^{d}$ that satisfies:
@@
$\left\{\begin{aligned} -\nabla\cdot\boldsymbol{\sigma}(\boldsymbol{x})&=\boldsymbol{f}(\boldsymbol{x})\,,\qquad&&\boldsymbol{x}\in~\Omega,\\ \boldsymbol{u}(\boldsymbol{x})&=\boldsymbol{0}\,,\qquad&&\boldsymbol{x}\in\Gamma_{D},\\ \boldsymbol{\sigma}(\boldsymbol{x})\cdot\boldsymbol{n}&=\boldsymbol{t}(\boldsymbol{x})\,,\qquad&&\boldsymbol{x}\in\Gamma_{N},\\ \boldsymbol{\sigma}(\boldsymbol{x})\cdot\boldsymbol{n}&=\boldsymbol{0}\,,\qquad&&\boldsymbol{x}\in\partial\Omega\setminus(\Gamma_{D}\cup\Gamma_{N}).\end{aligned}\right$. (1)
@@
where $\Gamma_{D}$ and $\Gamma_{N}$ denote the boundaries where Dirichlet (displacement) and Neumann (traction) conditions are imposed, respectively. $\boldsymbol{f}$ represents the body force, $\boldsymbol{t}$ represents the surface traction, and $\boldsymbol{n}$ is the unit outward normal vector to the boundary.
The governing equations of the system encompass two fundamental relationships. First is the geometric relation, which relates the strain tensor $\boldsymbol{\varepsilon}(\boldsymbol{x})$ to the displacement field $\boldsymbol{u}(\boldsymbol{x})$ (under the assumption of infinitesimal strain):
$\boldsymbol{\varepsilon}(\boldsymbol{x})=\frac{1}{2}\left(\nabla\boldsymbol{u}+(\nabla\boldsymbol{u})^{T}\right),$ (2)
@@
Second is the constitutive relation (Hooke’s law), linking the stress tensor $\boldsymbol{\sigma}(\boldsymbol{x})$ to the strain $\boldsymbol{\varepsilon}(\boldsymbol{x})$. Using Voigt notation, this relationship is expressed as:
$\boldsymbol{\sigma}(\boldsymbol{x})=\boldsymbol{C}(\boldsymbol{x})\boldsymbol{\varepsilon}(\boldsymbol{x}),$ (3)
here, $\boldsymbol{C}$ is the stiffness tensor, corresponding to a matrix in $\mathbb{R}^{3\times 3}$ for 2D cases and $\mathbb{R}^{6\times 6}$ for 3D cases, with its specific form determined by the material parameters.
@@
We discretize the domain $\Omega$ using a mesh (which may be structured or unstructured), denoting the triangular/tetrahedral mesh as $\mathcal{T}_{h}$ with a total of $N$ vertices. We employ the standard continuous Galerkin linear finite element method, defining the vector-valued Sobolev space $\boldsymbol{V}=[H^{1}(\Omega)]^{d}$. The corresponding discrete trial and test function spaces (which coincide due to the homogeneous Dirichlet boundary conditions) are defined as:
@@
$\boldsymbol{U}_{h}=\boldsymbol{V}_{h}=\left\{\boldsymbol{v}_{h}\in\boldsymbol{V}\mid\boldsymbol{v}_{h}|_{K}\in[\mathbb{P}_{1}(K)]^{d},\forall K\in\mathcal{T}_{h},\text{ and }\boldsymbol{v}_{h}|_{\Gamma_{D}}=\boldsymbol{0}\right\},$
@@
The discrete variational problem corresponding to (1) is: find $\boldsymbol{u}_{h}\in\boldsymbol{U}_{h}$ such that
$a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})=\ell(\boldsymbol{v}_{h})\quad\forall\boldsymbol{v}_{h}\in\boldsymbol{V}_{h}.$
The bilinear form and linear functional are given respectively by
$a(\boldsymbol{u}_{h},\boldsymbol{v}_{h})$ $=\int_{\Omega}\boldsymbol{\sigma}(\boldsymbol{u}_{h}):\boldsymbol{\varepsilon}(\boldsymbol{v}_{h})\,d\boldsymbol{x},$
$\ell(\boldsymbol{v}_{h})$ $=\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{v}_{h}\,d\boldsymbol{x}+\int_{\Gamma_{N}}\boldsymbol{t}\cdot\boldsymbol{v}_{h}\,ds.$
@@
After finite element discretization, we obtain the following system of linear algebraic equations
$\boldsymbol{A}\boldsymbol{u}=\boldsymbol{f},$ (4)
@@
where $\boldsymbol{A}\in\mathbb{R}^{dN\times dN}$ is the global stiffness matrix, and $\boldsymbol{u},\boldsymbol{f}\in\mathbb{R}^{dN}$ are the global displacement vector and load vector, respectively. We adopt a node-wise ordering of degrees of freedom (DOFs), which imparts a block structure to the system
$\boldsymbol{A}=\begin{pmatrix}\boldsymbol{A}_{11}&\cdots&\boldsymbol{A}_{1N}\\ \vdots&\ddots&\vdots\\ \boldsymbol{A}_{N1}&\cdots&\boldsymbol{A}_{NN}\end{pmatrix},\quad\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{u}_{N}\end{pmatrix},\quad\boldsymbol{f}=\begin{pmatrix}\boldsymbol{f}_{1}\\ \vdots\\ \boldsymbol{f}_{N}\end{pmatrix}$,
here, $\boldsymbol{A}_{ij}\in\mathbb{R}^{d\times d}$ represents the stiffness matrix blocks, while $\boldsymbol{u}_{i},\boldsymbol{f}_{i}\in\mathbb{R}^{d}$ correspond to the sub-vectors for the $i$-th node.
2.2 Graph Representation of Linear Systems

### v2

@@ Abstract
The efficient numerical solution of variable-coefficient linear elasticity equations on unstructured meshes presents a formidable challenge in computational mechanics. Convergence rates of classical iterative methods often stagnate due to material heterogeneity, strong anisotropy, and mesh irregularity. To address these limitations, this paper proposes a novel deep learning-based hybrid iterative method that integrates a weighted block Jacobi smoother with a graph neural network-enhanced global corrector. We introduce the adaptive graph Fourier neural solver, which employs a learnable coordinate transformation to construct a dynamic, problem-dependent spectral basis. This approach effectively overcomes the limitations of fixed Fourier bases in capturing the multi-scale features inherent in variable-coefficient media. Furthermore, to handle 3D and strongly anisotropic systems, we develop the multilevel adaptive graph Fourier neural solver, which executes hierarchical error correction across cascading frequency bandwidths. Rigorous theoretical analysis, grounded in the energy norm and Korn’s inequality, establishes mesh-independent convergence guarantees. Extensive numerical experiments on 2D and 3D elasticity problems demonstrate that the proposed method exhibits superior robustness and convergence rates compared to the classical smoothed aggregation algebraic multigrid method, serving effectively as both a standalone solver and a preconditioner for Krylov subspace methods.
@@ 1 Introduction
The linear elasticity model serves as the cornerstone of solid mechanics [atkin2013introduction, ting1996anisotropic, wang2008recent], characterizing the stress-strain behavior of deformable bodies subject to external loads. This framework is ubiquitous across engineering disciplines, including civil, aerospace, and mechanical engineering. Mathematically, the governing equations constitute a prototypical system of vector-valued partial differential equations (PDEs), distinguished by the inherent coupling between vector components. Driven by the escalating demand for high-fidelity simulations, solving these equations on complex geometries with heterogeneous materials imposes stringent requirements on the efficiency and robustness of computational algorithms [ciarlet2002finite, hughes2012finite].
@@
The finite element method (FEM) remains the predominant discretization technique due to its flexibility in handling complex geometries and multi-physics coupling [rannacher1997feed]. Discretization typically yields large-scale sparse linear systems. To solve such systems, classical iterative algorithms [saad2003iterative]—such as Jacobi, Gauss-Seidel, conjugate gradient, and generalized minimal residual (GMRES)—combined with preconditioning techniques like multigrid (MG) [trottenberg2000multigrid, li2025vertex] or domain decomposition methods, generally perform well. However, in regimes characterized by strong heterogeneity, pronounced anisotropy, or unstructured mesh topologies, these traditional methods often suffer from severe convergence degradation. Furthermore, constructing effective preconditioners for such scenarios is non-trivial and often necessitates heuristic parameter tuning. Consequently, achieving efficient solutions to complex systems while maintaining theoretical convergence guarantees remains a critical challenge in numerical analysis.
@@
Recently, deep learning has emerged as a transformative tool in scientific computing, offering novel perspectives on numerical method development. Research on solving physical problems, such as linear elasticity, generally follows two paradigms. The first utilizes deep neural networks to approximate the solution directly, exemplified by physics-informed neural networks, which embed PDE residuals into the loss function [wang2025novel]. The second focuses on neural operators, such as DeepONet [lu2021learning] and the Fourier neural operator (FNO) [lifourier], which learn the mapping from parameter space (e.g., material properties, external forces) to solution space, serving as surrogates for traditional solvers [taghikhani2025neural].
Despite their promise, methods relying entirely on neural surrogates often struggle to meet the accuracy and stability standards required for engineering applications. This limitation has prompted a shift towards integrating deep learning with traditional iterative methods, establishing the paradigm of “Neural Solvers.” Relevant research primarily falls into two categories: First, researchers aim to learn optimal parameters within traditional iterative schemes. These approaches preserve the mathematical structure of the original algorithms while utilizing neural networks to optimize specific components, thereby retaining theoretical properties. Examples include learning weights in Chebyshev iteration [WANG2026117013, cui2025neural], diagonal preconditioners in weighted Jacobi methods [moore2025graph], smoothers [katrutsa2020black, huang2022learning, chen2022meta] or prolongation/restriction operators in multigrid methods [greenfeld2019learning, luz2020learning, huang2024reducing], transmission conditions in domain decomposition [taghibakhshi2022learning], and initial guesses for Krylov subspace methods [arisaka2021gradient, NOVELLO2024112700]. Second, researchers attempt to replace part or all of the preconditioner’s functionality with neural networks. This includes developing deep learning-based hybrid iterative methods (DL-HIM), where simple iterative methods (e.g., Jacobi) eliminate high-frequency errors while neural networks correct low-frequency modes, forming a frequency-complementary strategy [zhang2024blending, cui2022fourier, cui2024hybrid, hu2025hybrid, hsieh2018learning, sun2025learning]. Another strategy involves neural operator preconditioning, where operators like FNO directly learn an approximate inverse of the system matrix for use within Krylov methods [dimola2025numerical, chen2025graph, trifonov2025efficient, xu2025neural].
@@
However, existing research on neural solvers has largely concentrated on scalar PDEs or structured grids. There is a paucity of studies addressing vector-valued PDEs on unstructured grids, which are prevalent in complex engineering analysis. From the perspective of numerical linear algebra, traditional iterative methods on unstructured grids face multiple hurdles, including convergence slowdowns due to mesh irregularity, ill-conditioning from high-contrast coefficients, and the difficulty of automating algorithm design. Therefore, constructing a solver framework that retains the numerical stability of traditional methods while leveraging the adaptability of neural networks on unstructured grids remains a significant open problem.
@@
To address these challenges, this paper proposes a hybrid solver that integrates weighted block Jacobi smoothing with a graph neural network (GNN)-enhanced Fourier error correction mechanism. GNNs are naturally suited for processing graph-structured data and can directly model the topological connectivity and nodal physical attributes of unstructured meshes, effectively preserving the coupling between displacement components. The weighted block Jacobi method rapidly attenuates certain error components, which, when combined with frequency-domain error correction, ensures effective treatment of the entire error spectrum. We establish a theoretical foundation using Korn’s inequality, ensuring robustness across different discretizations. We progressively develop the multilevel adaptive graph Fourier neural solver (ML-AG-FNS), incorporating learnable coordinate transformations and a hierarchical frequency correction mechanism to adaptively capture complex error modes induced by material heterogeneity and anisotropy.
The primary contributions of this work are as follows:
@@
    Building upon the Fourier neural solver (FNS) [cui2024hybrid], we propose a vertex-based, topology-aware iterative neural solver tailored for vector-valued PDEs on unstructured grids. By introducing a Fourier basis transformation network and integrating GNNs, this work substantially extends the applicability of the original FNS.
@@
    Through extensive numerical experiments on 2D and 3D linear elasticity problems involving isotropic heterogeneous and anisotropic materials, we demonstrate that the proposed solver exhibits superior robustness and efficiency, significantly outperforming the smoothed aggregation algebraic multigrid (SA-AMG) method [EnergyMinimization2011].
@@
    We provide a rigorous proof of unconditional convergence for the hybrid iterative scheme based on the energy norm. Leveraging Korn’s inequality, we derive mesh-independent bounds for the smoothing and correction steps, theoretically guaranteeing the stability and accuracy of the proposed learning-based solver.
@@
The remainder of this paper is organized as follows. Section 2 introduces the FEM of linear elasticity and the GNN-based implementation of weighted block Jacobi methods. Section 3 details the FNS framework and provides a convergence analysis. Section 4 describes the datasets. Section 5 presents the three progressively improved FNS architectures. Finally, Section 6 summarizes the contributions and discusses future directions.
2 Problem Formulation and Preliminaries
2.1 Linear Elasticity and Finite Element Discretization
@@
Consider the static linear elasticity equations defined on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d=2$ or $3$). We seek the displacement field $\bm{u}:\Omega\to\mathbb{R}^{d}$ governed by the following boundary value problem
@@
$\left\{\begin{aligned} -\nabla\cdot\bm{\sigma}(\bm{x})&=\bm{f}(\bm{x})\qquad&&\text{in }\Omega,\\ \bm{u}(\bm{x})&=\bm{0}\qquad&&\text{on }\Gamma_{D},\\ \bm{\sigma}(\bm{x})\cdot\bm{n}&=\bm{t}(\bm{x})\qquad&&\text{on }\Gamma_{N},\\ \bm{\sigma}(\bm{x})\cdot\bm{n}&=\bm{0}\qquad&&\text{on }\partial\Omega\setminus(\Gamma_{D}\cup\Gamma_{N}),\end{aligned}\right$. (1)
@@
where $\Gamma_{D}$ and $\Gamma_{N}$ denote the Dirichlet and Neumann boundaries, respectively. Here, $\bm{f}$ represents the body force, $\bm{t}$ is the surface traction, and $\bm{n}$ is the unit outward normal vector.
The system incorporates two fundamental physical laws. The first is the geometric relation between the linearized strain tensor $\bm{\varepsilon}$ and the displacement field $\bm{u}$, defined as
$\bm{\varepsilon}(\bm{u})=\frac{1}{2}\left(\nabla\bm{u}+(\nabla\bm{u})^{T}\right).$ (2)
@@
The second is the constitutive relation (Hooke’s Law) linking the stress tensor $\bm{\sigma}$ to the strain $\bm{\varepsilon}$. Using Voigt notation, this relationship is expressed as
$\bm{\sigma}(\bm{u})=\bm{C}\bm{\varepsilon}(\bm{u}),$ (3)
where $\bm{C}$ denotes the stiffness tensor. In 2D and 3D contexts, $\bm{C}$ is represented by matrices in $\mathbb{R}^{3\times 3}$ and $\mathbb{R}^{6\times 6}$, respectively, with entries determined by material properties.
@@
Let $\mathcal{T}_{h}$ be a partition of the domain $\Omega$ into triangular or tetrahedral elements, with a total of $N$ nodes. We employ the standard continuous Galerkin method with linear finite elements. The discrete function space, incorporating homogeneous Dirichlet boundary conditions, is defined as
@@
$\bm{V}_{h}=\left\{\bm{v}_{h}\in[H^{1}(\Omega)]^{d}\mathrel{\big|}\bm{v}_{h}|_{K}\in[\mathbb{P}_{1}(K)]^{d},\forall K\in\mathcal{T}_{h},\text{ and }\bm{v}_{h}|_{\Gamma_{D}}=\bm{0}\right\},$
@@
The corresponding discrete variational formulation seeks $\bm{u}_{h}\in\bm{V}_{h}$ such that
$a(\bm{u}_{h},\bm{v}_{h})=\ell(\bm{v}_{h})\quad\forall\bm{v}_{h}\in\bm{V}_{h},$
where the bilinear form $a(\cdot,\cdot)$ and the linear functional $\ell(\cdot)$ are given by
$a(\bm{u}_{h},\bm{v}_{h})$ $=\int_{\Omega}\bm{\sigma}(\bm{u}_{h}):\bm{\varepsilon}(\bm{v}_{h})\,d\bm{x},$ (4)
$\ell(\bm{v}_{h})$ $=\int_{\Omega}\bm{f}\cdot\bm{v}_{h}\,d\bm{x}+\int_{\Gamma_{N}}\bm{t}\cdot\bm{v}_{h}\,ds.$
@@
Finite element discretization yields the following system of linear algebraic equations
$\bm{A}\bm{u}=\bm{f},$ (5)
@@
where $\bm{A}\in\mathbb{R}^{dN\times dN}$ is the global stiffness matrix, while $\bm{u},\bm{f}\in\mathbb{R}^{dN}$ represent the global displacement and load vectors, respectively. By adopting a node-wise ordering of degrees of freedom (DOFs), the system exhibits a block structure
$\bm{A}=\begin{pmatrix}\bm{A}_{11}&\cdots&\bm{A}_{1N}\\ \vdots&\ddots&\vdots\\ \bm{A}_{N1}&\cdots&\bm{A}_{NN}\end{pmatrix},\quad\bm{u}=\begin{pmatrix}\bm{u}_{1}\\ \vdots\\ \bm{u}_{N}\end{pmatrix},\quad\bm{f}=\begin{pmatrix}\bm{f}_{1}\\ \vdots\\ \bm{f}_{N}\end{pmatrix}$,
where each block $\bm{A}_{ij}\in\mathbb{R}^{d\times d}$ couples the DOFs between nodes $i$ and $j$, and $\bm{u}_{i},\bm{f}_{i}\in\mathbb{R}^{d}$ are the sub-vectors associated with the $i$-th node.
2.2 Graph Representation and Message Passing Mechanism